# Desynchronization Index: A New Connectivity Approach for Exploring Epileptogenic Networks

### v2

@@ IV-B Performance Evaluation
The EI and DI algorithms are both implemented as standalone tools or combined according to the “and” and “or” Boolean functions. Practically, in the first case (named “EI and DI”), a channel is classified as epileptogenic only if both the algorithms mark it as positive. Instead, in the second case (named “EI or DI”), a channel is classified as epileptogenic if at least one of the algorithms classifies it as positive.
We first consider as ground truth the EZ as defined by the neurophysiologists; then, the analysis is repeated by considering a subset of the EZ channels, named Detectable Epileptogenic Zone (D-EZ). This latter includes all the EZ channels classified as positive by at least one of the algorithms (i.e., using the EI or DI approach). This approach makes it not possible to describe the algorithms’ performance in absolute terms but only to compare their results.
As the last step, we evaluate the performance of the algorithm in identifying the channels involved in the propagation of epileptic discharges. To this goal, we define a new detection class, named Detectable Propagation Zone (D-PZ), that includes all the channels classified as positive by at least one of the algorithms (i.e., using the “EI or DI” approach) and, afterward, labeled as propagator by the neurophysiologists. Although the EZ removal is the primary goal of epilepsy surgery, also the identification of propagator channels is extremely valuable for surgical planning. In the case that some epileptogenic sites cannot be removed (e.g., because this will endanger some cognitive functions of the patient), including the Propagation Zone (PZ) in the resection area may be the only solution for ensuring a seizure reduction.
@@
We observe that other works in this field consider as ground truth the area resected after epilepsy surgery or the target of thermocoagulation treatment. However, using these information may not enables the accurate definition of the EZ as, notably, epilepsy surgery is designed to remove a larger area than the sole EZ, to maximize the probability of seizure reduction in the patient. According to our framework, instead, the definition of the EZ is based on the clinical evaluation of the same SEEG data taken as input by our computational framework. This allows us to evaluate the EI and DI algorithms in fairly context, i.e., during the clinical stage in which it is expected to be used, against a pure human analysis.
For all the above reasons, relying on neurophysiological evaluation for the definition of the clinical ground truth represents one of the main strength of this manuscript. On the other hand, this prevents us from considering other SEEG datasets to further assess the accuracy of our algorithms, as pre-existent data are likely to be associated with labels different from those used in this work. In the future, repeating the analysis involving multiple centers, following a consistent methodology for SEEG analysis, will be paramount for the validation of the proposed computational framework.
V Results
@@
In this section, we present the results obtained when exploiting the EI and DI algorithms for detecting the EZ, considering the performance metrics given in Sec. IV-B. We remark that our computational framework is publicly available at the link: https://github.com/masonfed/desync_index. At the end of the section, we review in more detail two specific seizures among those analyzed, comparing the algorithm outcomes with clinical observations.
V-A Detection Results
@@
The SEEG signals are segmented in time windows lasting $T_{\text{window}}=1.00$ s each, considering a time shift of $\Delta t=0.25$ s. Given the epoch duration ($T_{\text{epoch}}=200$ s) a total of $797$ windows is analyzed for each seizure. For estimating the connectivity between different channels according to the PTE algorithm, we consider $\tau_{max}=0.10$ s as maximum propagation delay. Instead, for computing the energy ratio, the high- and low-frequency ranges are set to $B_{h}=[30.0,250.0]$ Hz and $B_{\ell}=[4.0,12.0]$ Hz, respectively. We observe that such a setting extends the portion of the spectrum processed by the EI algorithm, which, in its initial version, considered frequency lower than $97$ Hz because of the limited sampling rate [7]. In this work, we consider a wider frequency range to capture ripple phenomena that are notably characterized by oscillations up to $250$ Hz or even more [43]. On the other hand, this can lead to the inclusion of noisy information within our framework, leading to more false positive outcomes using either the EI and the DI algorithms.
In this work, we set the maximum number of epileptogenic channels to $M=15$, the tonicity interval to $\delta=5.00$ s (as done in [7]), the weight of the CUSUM chart to $\gamma=0.0$, and the exponential decay of the EWMA chart to $\alpha=0.1$. The tuning of $\gamma$ and $\alpha$ followed the recommendations provided in [44], and are kept to low values to capture the fast dynamics that characterized ictal phenomena. We set the detection threshold of the algorithms to $\eta=0.2$, which implies that we consider as part of the EZ only those channels presenting an epileptogenic value higher than $0.2$. Changing $\eta$ as well as $M$ affects the well-known trade-off between sensitivity and precision of detection algorithms. Finally, we set the starting time of the baseline period of the CUSUM chart to $t_{\text{base}}=t_{\text{start}}-50.0$ s. This implies that energy and connectivity statistics are estimating according to the signal evolution during the $50.0$ s preceding the seizure onset. As explained in Sec. IV, $t_{\text{start}}$ and $t_{\text{end}}$ depend on the specific seizure analyzed and, thus, are epoch-tailored.
(a) D-EZ.
(b) D-PZ.
Figure 2: Performance metrics in the D-EZ and D-PZ detection.
@@
In Fig. 1, we report the detection performance of EI, DI, and the algorithm combinations, considering the channels within the EZ as a classification target. We can appreciate that DI shows higher sensitivity than EI ($0.39$ and $0.32$ in the two cases, respectively) when they are used as a standalone tool, while their combination (EI or DI) increases the sensitivity of more than $50$% (with a final score of $0.54$). This denotes that two algorithms return different sets of results, i.e., the channels associated with a higher EI value are not the same associated with a higher DI value. Focusing on precision, the EI algorithm leads to the best results (with a score of $0.51$), while considering the “and” function to combine the algorithms leads to a performance gain of $30\%$. On the other hand, the DI algorithm leads to a higher number of false positives and is indeed associated with a precision of only $0.35$. In general, the best accuracy is obtained using the “EI or DI” approach, which returns a score of $0.96$.
In Fig. 2, we report the detection performance considering the channels of the D-EZ and D-PZ zones as a classification target. We recall that D-EZ and D-PZ are defined according to the output of the “EI or DI” system, including all the epileptogenic and propagation channels that any of the algorithms classify as positive. In both cases, the DI algorithm returns the best results in terms of sensitivity ($0.73$ and $0.64$, respectively), while the EI algorithm leads to higher precision. This confirms that the use of DI, enables a more holistic analysis of the overall ictal transition, including more channels among those suspected of being part of the EZ, but may produce many false positives in the case of complex network topologies. On the other hand, DI is not intended to directly plan epilepsy surgery but to support neurophysiological evaluation. In a practical scenario, neurophysiologists themselves can discern false and true positives for the algorithm outcomes, making the sensitivity the primary performance indicator for our framework.
V-B Concordant Example
@@
In the following, we better evaluate the operations of the two algorithms, analyzing two specific seizure events. We first consider patient $7$, whose SEEG implant (represented in Fig. 3(a)) explored the right temporo-parieto-occipital area, and includes $152$ different recording sites ($18$ electrodes). In Fig. 4(a), we report the bipolar representation of the $M=15$ channels associated with the strongest variation in the energy ratio during the 200 seconds preceding the propagation of epileptic discharges. Instead, in Fig. 4(b), we report the energy ratio values associated with each channel during the same period. In the two figures, the time is shifted according to the start of the ictal phase, so that $t_{\text{start}}=0$ s and $t_{\text{base}}=-50$ s. In particular, $t_{\text{base}}$ denotes the instant from which the baseline statistics of the energy ratio per channel were calculated, $t_{\text{base}}$ denotes the instant from which the algorithm starts detecting signal abnormalities, while the markers $t_{x}$ denote the activation times, i.e., the times at which the energy ratio $E_{x}$ diverges from its baseline.
@@
In Fig. 5, we represent the operation of the DI algorithm in the same scenario. Hence, Fig. 5(a) reports the $M=15$ channels showing the most significant connectivity changes, while Fig. 5(b)) reports the desynchronization level for each of the same channels. Comparing Fig. 5(b) with Fig. 4(b), we can observe that the energy ratio represents a less noisy metric than the desynchronization level. During the baseline, the energy ratio of each channel tends to zero, while, for $t_{\text{start}}>0$, most channels present abrupt energy changes, with a peak $E_{x}$ for channels Y4 and Y5. The desynchronization level is characterized by a much higher variability: many channels present bursts of reduced connectivity for $t_{\text{start}}<0$. On the other hand, Fig. 5(b) identifies desynchronization phenomena in cortical sites that, apparently, do not present variation in terms of fast oscillations. This is the case of channel V8, which starts desynchronizing from the rest of the network tens of seconds before the seizure onset but is not recognized as epileptogenic by the EI algorithm.
@@
Finally, in Fig. 6, we report the epileptogenicity levels assigned to each channel by the two algorithms, as well as the brain structure with which each channel is associated. In both cases, the output of the algorithms is normalized within the $[0,1]$ range, which means that epileptogenic values close to $1$ are associated with a greater influence in the diffusion of epileptic discharge. We can appreciate that the outputs are mostly concordant as both techniques identify the most epileptogenic channels within the posterior collateral sulcus and the fusiform gyrus. At the same time, DI is the sole tool that marks the channels within the the lingual gyrus, which was considered one of the trigger of the patient epilepsy.
V-C Discordant Example
@@
In the following, we consider patient $8$, whose SEEG implant (represented in Fig. 3(b)) explored the left fronto-temporo-parietal, area, and includes $116$ recording sites ($13$ electrodes). As done previously, we report the $M=15$ channels considered as most epileptogenic according to the two approaches, focusing first on the EI algorithm (Fig. 9(a)) and then on the DI algorithm (Fig. 9(b)). As before, the energy ratio associated with each channel presents low values during the baseline and then abruptly rises during the ictal phase. In particular, the channels associated with electrode I, crossing the temporal pole, maximize the energy ratio a few seconds after $t_{\text{start}}$.
In Fig. 9(b), we perform the same analysis for the desynchronization level. In this case, the network connectivity is less noisy than the seizure analyzed in Fig. 5 since the channels do not show prominent variations during the baseline period. This difference may be due to the different network sizes: the SEEG implant of patient $8$ includes $36$ more recording sites than that of patient $7$. In other words, in this case, the SEEG implant explores a much more focused epileptogenic network and, thus, leads to more stable connectivity dynamics.
Focus on the ictal phase ($t_{\text{start}}>0$), we can appreciate that DI identifies desynchronization phenomena in several structures that, instead, are not marked by epileptogenic in terms of energy ratio. As shown in Fig. 9(b), channels associated with electrodes B, C, and D (which all explore the cortical area neighboring the hyppocampus and parahyppocampus) are characterized by a significant drop in connectivity a few seconds after the start of the ictal phase. None of such channels (but B1) was included in those marked by EI, which can lead to the definition of a more restricted resection area during epilepsy surgery, with a lower chance of totally suppressing the mechanisms underlying seizure activation.
@@
Finally, Fig. 9 compares the epileptogenicity levels assigned to each channel by the algorithms, also reporting the cortical locations of the channels within the SEEG implants. In this case, the algorithm outputs is only partially concordant: EI marked the channels associated with the temporal pole as the most epileptogenic, while DI seems to identify the origin of the epileptic discharge in the hyppocampus. In particular, using EI, the most epileptogenic channel is I1, while the DI algorithm underscores channel C5, which is located within the hippocampal tail. This heterogeneous results denote how it is necessary to integrate multiple techniques, considering both energy-based and connectivity-based metrics, to improve the accuracy of EZ localization.
V-D Clinical Discussion
@@
Our results show that the DI algorithm can recognize the desynchronization phenomena that precede the generation of the epileptic discharge in electrodes that intersect or surround the EZ. This leads DI to have a partially concordant output with EI, which, notably, aims at quantifying the magnitude of the high-frequency energy associated with ictal activities. Contrary to EI and most state-of-the-art solutions, DI does not compute the signal energy over specific frequency bands, but evaluates the desynchronization level of each SEEG channel considering all its oscillatory components. This makes it possible to implement the algorithm in a patient-agnostic fashion, without the need for assumptions about the frequency ranges characterizing ictal phenomena. Refining DI to consider specific frequency range is also possible and may further improve the accuracy of our methodology, although this may constitute a biasing factor for the results and reduce the generalizability of the framework.
@@
When combined with EI, the proposed DI algorithm allows us to improve the accuracy in EZ definition and, used as a standalone tool, shows higher performance also in the detection of the cortical sites within the PZ. From a practical perspective, DI marks as epileptogenic the cortical sites that abruptly assume a static phase distribution in time and, thus, are not influenced by any other sites explored through the SEEG implant. Hence, the key idea behind our methodology is that epileptogenic signals present an independent behavior that cannot be inhibited by the other cortical structures. This is in line with the most recent findings in the field of EZ detection and confirms that the generation of epileptic seizures is allowed by a connectivity reduction in the overall network. Hence, during the inter-ictal phase, the regions surrounding the EZ may assume an inhibitory function that prevents the diffusion of epileptic discharges and, thus, constitute a key component of the epileptogenic network.
@@
In a clinical context, the DI algorithm is not intended to directly plan the resection area for epilepsy surgery, but constitutes a support for neurophysiologists to identify the EZ and cortical sites that contribute to the propagation of epileptic discharges. In particular, DI emphasizes phenomena that are not marked by other computational techniques, such as EI, and are not easy to identify by a pure visual analysis. Our results prove that the combined use of DI and EI can strongly increase sensitivity in detecting ictal phenomena, although marking as epileptogenic cortical sites that are not part of the EZ. From this perspective, the clinical judgment of neurophysiologists remains an essential element to discern false and true positive, counterbalancing the lower precision of our framework. On the other hand, we remark that discerning ictal and non-ictal channels in a limited set, as the one given by our framework is a rather low-cost task, while reviewing all the SEEG channels is highly expensive and may lead to fatal human errors.
@@
Finally, we observe that the proposed DI algorithm may also reflect the presence of ictal activities in cortical areas not directly explored by the SEEG implant. This may be useful in scenarios where the SEEG signal does not enable the clear visualization of the seizure generation and propagation. Hence, our technique may help neurophysiologists to discern whether the epileptogenic network is totally or only partially explored by the SEEG implant and, thus, the information obtained by the SEEG monitoring is sufficient to ensure, with high probability, the success of epilepsy surgery.
VI Conclusion

### v3

@@ 4.2 Performance Evaluation
In addition, we study the trade-off between sensitivity and precision by computing the Receiver Operating Characteristic (ROC) and the Area Under the ROC Curve (AUC). The AUC is a common performance indicator for diagnostic tools: in clinical settings, a value higher than $0.80$ is solely required for obtain a reliable solution [44]. Practically, for each patient, the ROC curve is obtained by varying the number $M$ of epileptogenic channels up to the size of the SEEG implant.
@@
The EI and DI algorithms are both implemented as standalone tools or combined according to the “and” and “or” Boolean functions. Practically, in the first case (named “EI and DI”), a channel is classified as epileptogenic only if both the algorithms mark it as positive. Instead, in the second case (named “EI or DI”), a channel is classified as epileptogenic if at least one of the algorithms classifies it as positive. Clearly, we expect the first configuration to provide higher sensitivity, and the second configuration to return less false positives.
@@
We observe that other works in this field consider as ground truth the area resected after epilepsy surgery or the target of thermocoagulation treatment. However, using this information does not ensure an accurate EZ definition as, notably, epilepsy surgery is designed to remove a larger area than the sole EZ. In this work, we defined the EZ according to the neurophysiological evaluation of the same SEEG records taken as input by our computational framework. This allows us to evaluate the EI and DI algorithms in fairly context, i.e., during the clinical stage in which it is expected to be used, against a pure human analysis.
5 Results
@@
In this section, we present the results obtained by exploiting the EI and DI algorithms for detecting the EZ, considering the performance metrics given in Sec. 4.2. Our computational framework is publicly available at the link: https://github.com/masonfed/desync_index. At the end of the section, we review in more detail two specific seizures among those analyzed, comparing the algorithm outcomes with clinical observations.
5.1 Detection Results
The SEEG signals are segmented in time windows lasting $T_{\text{window}}=1.00$ s each, considering a time shift of $\Delta t=0.25$ s. Given the epoch duration ($T_{\text{epoch}}=200$ s) a total of $797$ windows is analyzed for each seizure. For estimating the connectivity between different channels according to the PTE algorithm, we consider $\tau_{max}=0.10$ s as maximum propagation delay. Instead, for computing the energy ratio, the high and low frequency ranges are set to $B_{h}=[30.0,250.0]$ Hz and $B_{\ell}=[4.0,12.0]$ Hz, respectively. Such a setting extends the portion of the spectrum processed by the EI algorithm, which, in its initial version, considered frequencies lower than $97$ Hz [8]. Here, we consider a wider frequency range to capture ripple phenomena that are notably characterized by oscillations up to $250$ Hz or even more [45].
We set the tonicity interval to $\delta=5.00$ s (as done in [8]), the weight of the CUSUM chart to $\gamma=0.0$, and the exponential decay of the EWMA chart to $\alpha=1.0$, following the recommendations provided in [46]. We set the detection threshold of the algorithms to $\eta=0.0$, which means that no channels were discarded from the algorithm output and the number of epileptogenic channels is always equal to $M$. Finally, we set the baseline period of the CUSUM chart to $t_{\text{base}}=t_{\text{start}}-20.0$ s so that the energy and connectivity statistics of each channel are estimating according to the $20.0$ s preceding the onset of the seizure. As explained in Sec. 4, $t_{\text{start}}$ and $t_{\text{end}}$ depend on the specific seizure analyzed and, therefore, are epoch-tailored.
(a) $M=10\%$.
(b) $M=20\%$.
@@ 5.1 Detection Results
Figure 2: Area Under the ROC Curve (AUC) in the EZ detection.
@@
In Fig. 1(a), we report the detection performance of the different algorithms, considering the channels within the EZ as a classification target. In particular, we tune $M$ is such a way $10\%$ of the channels are considered epileptogenic. We can appreciate that DI has higher sensitivity than EI ($0.44$ and $0.40$, respectively), while the ”EI or DI” configuration increases the sensitivity more than $50$% with respect EI (with a final score of $0.60$). This denotes that two algorithms return different results, i.e., the channels associated with a higher EI value are not the same associated with a higher DI value. Focusing on precision, the DI algorithm still overcomes EI ($0.40$ vs $0.34$), while considering the “and” function leads to a performance gain of $20\%$. Overall, the best accuracy is obtained using the “EI or DI” approach, which returns a score of $0.95$, followed by the DI algorithm at second place.
To improve sensitivity, we can increase the number $M$ of channels returned by the algorithms. When we set $M=20\%$, the sensitivity of the “EI or DI” approach reaches a value of $0.77$ while the accuracy decreases to $0.94$ due to the higher number of false positives (Fig. 1(b)). On the other hand, we highlight that the main advantage of our technique is to build first sketch of EZ, allowing neurophysiologists to look at a limited number of channels relative to the original network. This is strongly beneficial in the case of extensive SEEG implants (with hundreds of channels), where understanding the time-varying dynamics that lead to seizures is extremely difficult with a pure visual analysis.
To comprehensively analyze the trade-off between sensitivity and specificity, we vary the value of $M$ from $1\%$ to $100\%$ and compute the AUC associated with each technique. In particular, $M=100\%$ represents the extreme scenario in which all SEEG channels are marked as epileptogenic and neurophysiologists must look at the SEEG records from scratch. In this perspective, we may consider $M$ as a marker of the effort that neurophysiologists must exert and, therefore, the effectiveness of our framework.
Fig. 2 denotes how the ”EI or DI” approach leads to the best detection performance ($AUC=0.87$), followed by the DI algorithm ($AUC=0.80$) and the EI algorithm ($AUC=0.74$) at third place. These results confirm that the combined use of EI and DI enables a more holistic analysis of the overall ictal transition, allowing us to identify most of the channels that are part of the EZ. On the other hand, to ensure a sensitivity close to $0.8$, we must still look at more than $M=20\%$ of the SEEG channels (as we observed in Fig. 1(b)). We expect that integration of the proposed tools with other quantitative features, for example, taking into account the space position of recording sites within the implant, can further enhance the detection performance. In a clinical setting, this will make it possible to further reduce the neurophysiologist effort, i.e., the number $M$ of channels to be individually evaluated.
@@
5.2 Concordant Example
In the following, we better evaluate the operations of the two algorithms, analyzing two specific seizure events. We first consider patient $7$, whose SEEG implant (represented in Fig. 3(a)) explored the right temporo-parieto-occipital area, and includes $152$ different recording sites ($18$ electrodes). In Fig. 4(a), we report the bipolar representation of the $M=10$ channels associated with the strongest variation in the energy ratio during the 50 seconds preceding the propagation of epileptic discharges. Instead, in Fig. 4(b), we report the energy ratio values associated with each channel.
In the two figures, the time is shifted according to the start of the ictal phase, so that $t_{\text{start}}=0$ s and $t_{\text{base}}=-20$ s. In particular, $t_{\text{base}}$ denotes the instant from which the baseline statistics of the energy ratio per channel were calculated, $t_{\text{base}}$ denotes the instant from which the algorithm starts detecting signal abnormalities, while the markers $t_{x}$ denote the activation times, i.e., the times at which the energy ratio $E_{x}$ diverges from its baseline.
@@ 5.2 Concordant Example
In Fig. 5, we represent the operation of the DI algorithm in the same scenario. Hence, Fig. 5(a) reports the $M=10$ channels showing the most significant connectivity changes, while Fig. 5(b)) reports the desynchronization level for each of the same channels. Comparing Fig. 5(b) with Fig. 4(b), we can observe that the energy ratio represents a less noisy metric than the desynchronization level. During the baseline, the energy ratio of each channel tends to zero, while, for $t_{\text{start}}>0$, most channels present abrupt energy changes, with a peak $E_{x}$ for channels Y4 and Y5. The desynchronization level is characterized by higher variability, and many channels present bursts of reduced connectivity for $t_{\text{start}}<0$. In particular, channels Y4 and Y5 present abnormal connectivity patterns even during the baseline.
@@
Hence, in Fig. 6, we report the epileptogenicity levels assigned to each channel by the two algorithms, as well as the brain structure with which each channel is associated. In both cases, the output of the algorithms is normalized within the $[0,1]$ range so that values close to $1$ are associated with a greater influence on the diffusion of epileptic discharge. We can appreciate that the outputs are mostly concordant as both techniques identify the most epileptogenic channels within the posterior collateral sulcus and the fusiform gyrus, which were considered triggers of the patient’s seizures.
5.3 Discordant Example
@@
In the following, we consider patient $8$, whose SEEG implant (represented in Fig. 3(b)) explored the left fronto-temporo-parietal, area, and includes $116$ recording sites ($13$ electrodes). We report the $M=10$ channels considered the most epileptogenic according to the two approaches, focusing first on the EI algorithm (Fig. 7) and then on the DI algorithm (Fig. 8). As before, the energy ratio associated with each channel presents low values during the baseline and then increases abruptly during the ictal phase. In particular, the channels associated with electrode I, which crosses the temporal pole, maximize the energy ratio a few seconds after $t_{\text{start}}$.
In Fig. 8, we perform the same analysis for the desynchronization level. In this case, the network connectivity is slightly less noisy than the seizure analyzed in Fig. 5 since the channels do not show prominent variations during the baseline period. This difference may be due to the different network sizes: the SEEG implant of patient $8$ includes $36$ more recording sites than that of patient $7$. In other words, the SEEG implant explores a much more focused epileptogenic network and, thus, leads to a more stable connectivity dynamics.
Focusing on the ictal phase ($t_{\text{start}}>0$), we can appreciate that DI identifies desynchronization phenomena in several structures that, instead, are not marked by epileptogenic in terms of energy ratio. As shown in Fig. 8, channels associated with electrodes B, D, and E (which explore the cortical area adjacent the hyppocampus and collateral sulcus) are characterized by a significant decrease in connectivity a few seconds after the start of the ictal phase. None of these channels (but B1) was included in those marked by EI, which can lead to the definition of a more restricted resection area during epilepsy surgery, with a lower chance of totally suppressing the mechanisms underlying seizure activation.
@@
Fig. 9 compares the epileptogenicity levels assigned to each channel by the algorithms, also reporting the cortical locations of the channels within the SEEG implants. In this case, the output of the algorithm is only partially concordant: EI marked the channels associated with the temporal pole as the most epileptogenic, while DI seems to identify the origin of epileptic discharges in the limbic system. Using EI, the most epileptogenic channel is I1, while the algorithm DI places more emphasis on channel B1, B2, E4 and E5. These heterogeneous results denote the importance of integrating multiple techniques, considering both energy-based and connectivity-based metrics, to improve accuracy in the EZ localization.
5.4 Clinical Discussion
@@
Our results show that the DI algorithm can recognize the desynchronization phenomena that precede the propagation of epileptic discharge in electrodes that intersect or surround the EZ. This leads DI to have a partially concordant output with EI, which, notably, aims to quantify the magnitude of the high-frequency energy associated with ictal activities. Contrary to EI and most state-of-the-art solutions, DI does not compute the signal energy over specific frequency bands, but evaluates the desynchronization level of each SEEG channel considering all its oscillatory components. This makes it possible to implement the algorithm in a patient-agnostic fashion, without the need for assumptions about the frequency ranges characterizing ictal phenomena. Tuning DI over a specific frequency range is also possible and may further improve the accuracy of our methodology. At the same time, adopting a frequency-specific solution may constitute a bias and reduce the generalizability of the framework.
@@
When combined with EI, the DI algorithm allows us to greatly improve the accuracy in EZ localization. From a practical perspective, DI marks as epileptogenic the cortical sites that assume a static phase distribution over time and, thus, are not influenced by any other sites explored through the SEEG implant. The key idea behind our methodology is that epileptogenic signals present an independent behavior that cannot be inhibited by the other cortical structures. This is in line with the most recent findings in the field of EZ detection and confirms that the generation of epileptic seizures is allowed by a sudden reduction in network connectivity [47, 33, 48]. In this vision, the regions surrounding the EZ assume an inhibitory function that prevent epileptic discharges during the interictal phase.
@@
In a clinical context, DI is not intended to directly plan the resection area for epilepsy surgery, but to help neurophysiologists identify the cortical sites that contribute to the propagation of epileptic discharges. In particular, DI emphasizes phenomena that are not marked by other computational techniques, such as EI, and are not easy to identify by a pure visual analysis. The combined use of DI and EI can strongly increase sensitivity in the detection of epileptogenic sites and can be used to reduce the number of SEEG channels that must be reviewed in practical scenarios. At the same time, the clinical judgment of neurophysiologists remains an essential element to discern false and true positive, counterbalancing the lower precision of our framework. In this regard, we remark that discerning ictal and non-ictal channels in a limited set, as the one given by our framework is a rather low-cost task, while reviewing all the SEEG channels is highly expensive and may lead to fatal human errors.
@@
Finally, we observe that the DI algorithm may also reflect the presence of ictal activities in cortical areas not directly explored by the SEEG implant. This may be useful in scenarios where the SEEG signal does not provide a clear visualization of the ictal transition and the propagation of the seizures. Therefore, our framework can help neurophysiologists discern whether the epileptogenic network is totally or partially explored by the SEEG implant and, thus, the information obtained from the data is sufficient to ensure, with high probability, the success of epilepsy surgery.
6 Conclusion